A sample of a gas occupies a volume of 1100 mL at 78ºC and 0.060 atm. What would be the new volume if the temperature decreased to 44ºC and the pressure increased to 1.31 atm?
1
Expert's answer
2017-12-22T04:39:31-0500
p1 = 0,06 atm = 6079,5Pa
V1 = 1100ml = 0,0011m3
T1 = 780C = 351 K
p2 = 1,31 atm = 132735,8Pa
T2 = 440C = 317 K
V2 - ?
p1* V1 = m/M*R* T1
p2* V2 = m/M*R* T2
p1* V1 p2* V2 = T1 T2
V2 = p1* V1*T2 p2* T1 =6079,5Pa* 0,0011m3*317 K 132735,8Pa* 351 K = 4,5*10^-5 m3